80 students in the 5th grade voted for their favorite animal after a trip to the zoo. The elephants received 3/5 of the votes. How many votes did the elephant receive.
Answer by stanbon(75887)

80 students in the 5th grade voted for their favorite animal after a trip to the zoo. The elephants received 3/5 of the votes. How many votes did the elephant receive.
----------
Equation:
# of elephant votes = (3/5)80 = 3*16 = 48
====================
Cheers,
Stan H.
====================